# RLVR-World: Training World Models with Reinforcement Learning

### v1

@@ Appendix D Prompt Examples
[p] Text Game: Action Rule Example put: Description: put an object into a target container Rules: 1. The target must be a container (Container) 2. The target container must be open 3. The object must be in the inventory 4. The object must be moveable (isMoveable)
@@
[p] Text Game: Object Rule Example Object: Container Description: Abstract class for things that can be considered ’containers’ (e.g. a drawer, a box, a table, a shelf, etc.) Properties: - A Container is a container. - A Container could be opened (e.g., e.g. a drawer, a door, a box, etc.), or is it always "open" (e.g. a table, a shelf, etc.). - A Container has a property indicating if it is opened. - A Container has a property indicating the prefix to use when referring to the container (e.g. "in the drawer", "on the table", etc.). By default, the prefix is "in"
@@
[p] Text Game: Score Rule Example The player wins the game by getting all dishes clean. The player gets one point for each dish that is cleaned. The player loses one point for each dish that is made dirty.
@@
[p] Web Page: Web State Prediction Prompt Example You are an intelligent agent that predicts the next state from a given current action using your own logical reasoning. You will be given web-based tasks. Here’s the information you’ll have: - The user’s objective: This is the task you’re trying to complete. - The current web page’s accessibility tree: A simplified representation of the webpage, providing key information. - The current web page’s URL: This is the page you’re currently navigating. - The previous action: The action you just performed. It may help you track your progress. - The current action: The action that you will perform next to achieve the user’s objective in the current accessibility tree. The format of previous and current actions can fall into several categories: Page Operation Actions: click [id]: Click an element with a specific id. type [id] [content]: Type the content into the field with the specified id. By default, the "Enter" key is pressed after typing unless [0] is appended. hover [id]: Hover over an element. press [key_comb]: Simulate pressing a keyboard combination (e.g., Ctrl+v). scroll [down] or scroll [up]: Scroll the page. Tab Management Actions: new_tab: Open a new browser tab. tab_focus [tab_index]: Switch focus to a specific tab. close_tab: Close the currently active tab. URL Navigation Actions: goto [url]: Navigate to a specific URL. go_back: Go to the previous page. go_forward: Go forward (if a previous go_back was executed). Completion Action: stop [answer]: Use this when you believe the task is complete. Provide the final answer in brackets if applicable. To succeed, it’s essential to understand how the current action impacts the next state of the webpage. You must verify whether the current action successfully produces the intended effect. Reasoning Guidelines: 1. Generate your answer using logical REASONING. 2. Directly predict the next state based on the current action. 3. Format your prediction exactly as described below. Format Description: The output is divided into three sections: New items: Deleted items: Updated items: Each section should be separated by an empty line. If a section has no content, write None.
@@
[…]
@@
[p] Web Agent: Next-State Change Summarization Prompt You are an intelligent agent that predicts the next state based on the current action using your own logical reasoning. You will be given a web-based task. You will have access to the following information: • The user’s objective: The task you’re trying to complete. • The current observation: A simplified representation of the current webpage. • The current URL: The webpage you’re currently on. • Previous actions: The action(s) performed just prior to the current one. • Current action: The action being evaluated for its effect on the webpage. • Key changes in next state observation: A summary of differences between the current and next state observations. Action types may include: Page Operation Actions: click [id]: Click an element with the given ID. type [id] [content]: Type content into a field with the given ID (append [0] to suppress Enter). hover [id]: Hover over an element. press [key_comb]: Simulate key combination presses (e.g., Ctrl+v). scroll [down] or scroll [up]: Scroll the page. Tab Management Actions: new_tab, tab_focus [index], close_tab URL Navigation Actions: goto [url], go_back, go_forward Completion Action: stop [answer]: Use this when the task is complete. Provide the answer in the brackets if required. Your task: Predict the next state with proper reasoning. Follow these rules: 1. Begin your answer with: [Rationale] Let’s think step by step... 2. Your reasoning must mention the key changes in the next state observation. 3. Then, describe the next state based on those key changes. 4. Output the prediction in this format: [Next State] The expected effect is that ... Input fields: URL: {url} OBJECTIVE: {objective} PREVIOUS ACTION: {prev_action} CURRENT OBSERVATION: {cur_observation} CURRENT ACTION: {cur_action} KEY CHANGES IN NEXT STATE OBSERVATION: {tao}
@@
[p] Web Agent: Web State Evaluation Prompt Example You are an expert in evaluating and guiding a web navigation agent. Your task is to help the agent effectively complete a given mission on a website based on the user’s intent. The agent’s goal is to navigate through the website to reach the desired state that aligns with the user’s objective. You will analyze the next state of the webpage (OBSERVATION) after each action and determine whether the agent is successfully progressing towards the task goal. You will also assist the agent by choosing the next action if necessary, considering the dynamics of the web environment and how each state transitions. Key Points: 1. Understand the Intent: Identify the user’s goal (e.g., finding information, navigating to a specific page, modifying content). Ensure the next state of the webpage aligns with achieving that goal based on the current state and user’s intent. 2. Evaluate the Next State: Assess how the next state contributes to reaching the intended goal. If the next state moves the agent closer to the user’s goal, evaluate it positively. If the next state does not progress towards the goal or leads to an error, suggest alternative actions. 3. State Guidance: If the agent is on the right track but hasn’t completed the task yet, recommend further actions. Guide the agent toward a state that reflects clear progress towards the goal. 4. Types of Tasks: Information Seeking: The next state must provide specific information (e.g., product price, reviews). Site Navigation: The next state must reflect navigation to the exact page or item. Content Modification: The next state must show that content has been successfully modified. General Task: The final state should reflect that the task is completed. Only issue a stop action when the objective is met. 5. Common Pitfalls: Avoid corrupted input from repeated typing actions. Ensure the agent navigates to the specific item or content, not just to a general page. Output Format with Likert Scale: Each next state will be evaluated on a fine-grained scale from 1 to 5: 1: The next state is a failure or leads away from the task. 2: The next state does not contribute meaningfully. 3: The next state is neutral; the agent is maintaining position. 4: The next state is helpful; progress is being made. 5: The next state is optimal; directly aligns with the task goal. Response Instructions: Write a rationale providing a detailed analysis of the next state and justify your chosen score. Output Format: [Rationale]: <your thought> [Score]: <1--5>

### v2

@@ Appendix D Prompt Examples
Text Game: Action Rule Example put: Description: put an object into a target container Rules: 1. The target must be a container (Container) 2. The target container must be open 3. The object must be in the inventory 4. The object must be moveable (isMoveable)
@@
Text Game: Object Rule Example Object: Container Description: Abstract class for things that can be considered ’containers’ (e.g. a drawer, a box, a table, a shelf, etc.) Properties: - A Container is a container. - A Container could be opened (e.g., e.g. a drawer, a door, a box, etc.), or is it always "open" (e.g. a table, a shelf, etc.). - A Container has a property indicating if it is opened. - A Container has a property indicating the prefix to use when referring to the container (e.g. "in the drawer", "on the table", etc.). By default, the prefix is "in"
@@
Text Game: Score Rule Example The player wins the game by getting all dishes clean. The player gets one point for each dish that is cleaned. The player loses one point for each dish that is made dirty.
@@
Web Page: Web State Prediction Prompt Example You are an intelligent agent that predicts the next state from a given current action using your own logical reasoning. You will be given web-based tasks. Here’s the information you’$l$ have: - The user’s objective: This is the task you’re trying to complete. - The current web page’s accessibility tree: A simplified representation of the webpage, providing key information. - The current web page’s URL: This is the page you’re currently navigating. - The previous action: The action you just performed. It may help you track your progress. - The current action: The action that you will perform next to achieve the user’s objective in the current accessibility tree. The format of previous and current actions can fall into several categories: Page Operation Actions: click [id]: Click an element with a specific id. type [id] [content]: Type the content into the field with the specified id. By default, the "Enter" key is pressed after typing unless [0] is appended. hover [id]: Hover over an element. press [key_comb]: Simulate pressing a keyboard combination (e.g., Ctrl+v). scroll [down] or scroll [up]: Scroll the page. Tab Management Actions: new_tab: Open a new browser tab. tab_focus [tab_index]: Switch focus to a specific tab. close_tab: Close the currently active tab. URL Navigation Actions: goto [url]: Navigate to a specific URL. go_back: Go to the previous page. go_forward: Go forward (if a previous go_back was executed). Completion Action: stop [answer]: Use this when you believe the task is complete. Provide the final answer in brackets if applicable. To succeed, it’s essential to understand how the current action impacts the next state of the webpage. You must verify whether the current action successfully produces the intended effect. Reasoning Guidelines: 1. Generate your answer using logical REASONING. 2. Directly predict the next state based on the current action. 3. Format your prediction exactly as described below. Format Description: The output is divided into three sections: New items: Deleted items: Updated items: Each section should be separated by an empty line. If a section has no content, write None.
@@
Web Agent: Top-10 Salient Change Extraction Prompt Summarize the key changes in the web page based on the following information. New items: {new_items} Updated items: {updated_items} Deleted items: {deleted_items} When summarizing, follow this output format: 1. [First key change] 2. [Second key change] 3. [Third key change] $\vdots$ 10. [Tenth key change]
@@
Web Agent: Next-State Change Summarization Prompt You are an intelligent agent that predicts the next state based on the current action using your own logical reasoning. You will be given a web-based task. You will have access to the following information: • The user’s objective: The task you’re trying to complete. • The current observation: A simplified representation of the current webpage. • The current URL: The webpage you’re currently on. • Previous actions: The action(s) performed just prior to the current one. • Current action: The action being evaluated for its effect on the webpage. • Key changes in next state observation: A summary of differences between the current and next state observations. Action types may include: Page Operation Actions: click [id]: Click an element with the given ID. type [id] [content]: Type content into a field with the given ID (append [0] to suppress Enter). hover [id]: Hover over an element. press [key_comb]: Simulate key combination presses (e.g., Ctrl+v). scroll [down] or scroll [up]: Scroll the page. Tab Management Actions: new_tab, tab_focus [index], close_tab URL Navigation Actions: goto [url], go_back, go_forward Completion Action: stop [answer]: Use this when the task is complete. Provide the answer in the brackets if required. Your task: Predict the next state with proper reasoning. Follow these rules: 1. Begin your answer with: [Rationale] Let’s think step by step... 2. Your reasoning must mention the key changes in the next state observation. 3. Then, describe the next state based on those key changes. 4. Output the prediction in this format: [Next State] The expected effect is that ... Input fields: URL: {url} OBJECTIVE: {objective} PREVIOUS ACTION: {prev_action} CURRENT OBSERVATION: {cur_observation} CURRENT ACTION: {cur_action} KEY CHANGES IN NEXT STATE OBSERVATION: {tao}
@@
Web Agent: Web State Evaluation Prompt Example You are an expert in evaluating and guiding a web navigation agent. Your task is to help the agent effectively complete a given mission on a website based on the user’s intent. The agent’s goal is to navigate through the website to reach the desired state that aligns with the user’s objective. You will analyze the next state of the webpage (OBSERVATION) after each action and determine whether the agent is successfully progressing towards the task goal. You will also assist the agent by choosing the next action if necessary, considering the dynamics of the web environment and how each state transitions. Key Points: 1. Understand the Intent: Identify the user’s goal (e.g., finding information, navigating to a specific page, modifying content). Ensure the next state of the webpage aligns with achieving that goal based on the current state and user’s intent. 2. Evaluate the Next State: Assess how the next state contributes to reaching the intended goal. If the next state moves the agent closer to the user’s goal, evaluate it positively. If the next state does not progress towards the goal or leads to an error, suggest alternative actions. 3. State Guidance: If the agent is on the right track but hasn’t completed the task yet, recommend further actions. Guide the agent toward a state that reflects clear progress towards the goal. 4. Types of Tasks: Information Seeking: The next state must provide specific information (e.g., product price, reviews). Site Navigation: The next state must reflect navigation to the exact page or item. Content Modification: The next state must show that content has been successfully modified. General Task: The final state should reflect that the task is completed. Only issue a stop action when the objective is met. 5. Common Pitfalls: Avoid corrupted input from repeated typing actions. Ensure the agent navigates to the specific item or content, not just to a general page. Output Format with Likert Scale: Each next state will be evaluated on a fine-grained scale from 1 to 5: 1: The next state is a failure or leads away from the task. 2: The next state does not contribute meaningfully. 3: The next state is neutral; the agent is maintaining position. 4: The next state is helpful; progress is being made. 5: The next state is optimal; directly aligns with the task goal. Response Instructions: Write a rationale providing a detailed analysis of the next state and justify your chosen score. Output Format: [Rationale]: <your thought> [Score]: <1--5>